OpenStudy (kmullis6):

PLEASE HELP!!!!! Two of your friends, Matt and Karen, both run to you to settle a dispute. They were working on a math problem and got different answers. Wisely, you decide to look at their work & find the source of confusion. Matt 6 – 4(3 – 5)2 + 30 ÷ 5 6 – 4(–2)2 + 30 ÷ 5 6 – 4(4) + 30 ÷ 5 6 – 16 + 30 ÷ 5 −10+ 30 ÷ 5 20÷ 5 4 Karen 6 – 4(3 – 5)2 + 30 ÷ 5 6 – 4(–2)2 + 30 ÷ 5 6 – 4(−4) + 30 ÷ 5 6 + 16 + 30 ÷ 5 6 + 16 + 6 22+ 6 28

OpenStudy (kmullis6):

Explain to Matt and Karen who, if any, is correct and identify errors that you find. Provide the correct manner to fix those solutions and identify the correct answer. Use complete sentences.

OpenStudy (kmullis6):

@tHe_FiZiCx99

OpenStudy (anonymous):

Matt is wrong and Karen is correct

OpenStudy (anonymous):

The problem is 6-4(3-5)2+30/5 So following the order of operations you would do what is in the parenthesis first 3-5=-2 Now your math problem looks like 6-4(-2)2+30/5 You still have a parenthesis so now you do -2*2=-4 Now your math problem looks like 6-4*-4+30/5 Next step is multiplication so .. -4*-4=16 And now your math problem looks like 6+16+30/5 next step following the order of operations would be division so 30/5= 6 Now your math problem looks like 6+16+6 And from there you just add the numbers from left to right And your answer is 28 so Karen is definitely correct and all of her steps are in order !

OpenStudy (anonymous):

Matt's first mistake was when he did -2*2 and got positive 4 it's definitely negative 4 so that would make his problem incorrect from there and then he added before he divided your supposed to follow the order of operations when doing problems like this PEMDAS: Parenthesis Exponents Multiply Divide Add Subtract

OpenStudy (anonymous):

Hope that I've helped you as much as possible

OpenStudy (kmullis6):

Thank you so much. This helped me understand so much! (: @Shay17

OpenStudy (anonymous):

Lol You are very welcome let me know if you need any more help :)

OpenStudy (anonymous):

Thanks for the medal and the fan

OpenStudy (kmullis6):

I have a question about something tho.. When you said Matt was wrong because he did -2*2 and it was supposed to be -4. Two negatives make a positive. @Shay17

OpenStudy (kmullis6):

@Shay17

OpenStudy (anonymous):

Yea but you don't have two negatives you have a negative and a positive NEGATIVE 2 multiplied by POSITIVE 2 equals NEGATIVE 4

OpenStudy (kmullis6):

Omg sorry i read it wrong, i've been dealing with exponents A LOT lately, and i guess i read over that my bad lol :P @Shay17

OpenStudy (anonymous):

It's cool np

OpenStudy (kmullis6):

thanks for the fan btw (: @Shay17